OpenStudy (zubhanwc3):

a biker is slowing down according to the following table of velocities time: 0 1 2 3 4 velocity:25 20 16 12 10 use the trapezoidal rule to approximate the distance traveled in feet, during these 4 seconds velocity is in feet/sec a)64.1 b)65.5 c)66.5 d)67.4 e)83.0

OpenStudy (zubhanwc3):

im guessing its e

OpenStudy (anonymous):

The average speed between t=0 sec and t=1 sec is approximately (25+20)/2 ft/sec, using the trapezoid approximation. Therefore the distance traveled during this time is (25+20)/2 ft. Using the same approximations for the other time intervals, we estimate the distance traveled (in feet) as (25+20)/2 + (20+16)/2 + (16+12)/2 + (12+10)/2 = 25/2 + 20 + 16 + 12 + 10/2 = 65.5.

OpenStudy (zubhanwc3):

i was right, its e, u added it wrong, but you gave me the right concept.

OpenStudy (zubhanwc3):

nvm, u were right :/